The plane lx+my+nz=0 moves in such a way that it's intersection with the planes ax+by+cz+d=0 , a'x+b'y+c'z+d'=0 are perpendicular. Show that the normal to the plane through the origin describes, in general, a cone of second degree , and find its equation . Examine the case when aa'+bb'+cc'=0 , that is , when the two given planes are perpendicular.


1
Expert's answer
2021-03-03T07:02:17-0500

The plane "lx+my+nz=0" moves in such a way that it's intersection with the planes "ax+by+cz+d=0" , "a'x+b'y+c'z+d'=0" are perpendicular


The intersection with first plane gives the following line,


"\dfrac{x}{cm - bn} = \dfrac{y}{an - cl} = \dfrac{z}{bl - am}"


Similarly the intersection with the second line is :


"\dfrac{x}{c'm - b'n} = \dfrac{y}{a'n - c'l} = \dfrac{z}{b'l - a'm}"


For these lines to be perpendicular the direction ratio's inner product should be 0 , i.e.


"(cm - bn)(c'm - b'n) + (an - cl)(a'n - c'l) + (bl - am)(b'l - a'm) = 0"


Hence, this is equation of cone of second degree.